OpenStudy (anonymous):

what is the solution set of the inequality |12-3x| greater than or equal to 21

OpenStudy (john_es):

The solution is, \[x∈(−∞,−3]∪[11,∞)\] Split it in two, For x<4 \[-12+3x\geq 21\] For x>4 \[12-3x\geq 21\] And solve it simultaneosly,

OpenStudy (john_es):

You will find that, \[x\geq 33/3\Rightarrow x\geq 11\]\[-3x\geq 9\Rightarrow x\leq -3\]

OpenStudy (john_es):

I was wrong in the first post I deleted, the solution was easy. I hope it helps.

OpenStudy (anonymous):

thanks

OpenStudy (john_es):

You're welcome.